OpenStudy (anonymous):

ok, i need to simplyfy this and show all working: 3^3 / sqrt3^5

OpenStudy (anonymous):

\[3^{3}\div \sqrt{3^{5}}\]

OpenStudy (nowhereman):

You should use \[\sqrt{x} = x^{\frac{1}{2}}\]

OpenStudy (anonymous):

how

OpenStudy (nowhereman):

use power rules, and \[\frac{x}{y} = x\cdot y^{-1}\]

OpenStudy (nowhereman):

You do have the same bases there, so all you need no know then is what \[x^a \cdot x^b\] is

OpenStudy (anonymous):

oh, so it will be like \[3^{3-(1/5)}\] right?

OpenStudy (nowhereman):

nearly, but: \[\left( x^a \right) ^b = x^{a\cdot b} \]

OpenStudy (anonymous):

\[\sqrt{3^{5}} = 3^{1/5} . so you get 3^{3}/ \]

OpenStudy (anonymous):

you get 3^3 / 3^1/5

OpenStudy (anonymous):

sqrt3

OpenStudy (anonymous):

hence that will be \[3^{3-1/5}\]

OpenStudy (nowhereman):

that is exactly the mistake! \[ \sqrt{3^5} = \left( 3^5 \right)^{\frac{1}{2}}\]

OpenStudy (anonymous):

oh no i think i get it wrong. that should be 3^5/2

OpenStudy (anonymous):

hence, its \[3^{3-}\]

OpenStudy (anonymous):

My answer is true SQRT(3)

OpenStudy (anonymous):

its 3^3-5/2 = 3^1/2

OpenStudy (nowhereman):

And just for completeness, \[x^{\frac{1}{5}} = \sqrt[5]{x} \]

OpenStudy (anonymous):

the answer is \[\sqrt{3}\] right?

OpenStudy (anonymous):

duc, how do you get sqrt3?

OpenStudy (anonymous):

I know the answere, but I dont know how to get to the answer

OpenStudy (anonymous):

\[3^{3-5/2}\]

OpenStudy (anonymous):

ok, guys, i appreciate you helping me, i just became your fan

OpenStudy (anonymous):

First , You 3^5=3*3^4, so SQRT(3^5)=SQRT(3)*3^2,===>(3^3)/(SQRT(3)*3^2)=3/SQRT(3)=SQRT(3) It may be TRUE

OpenStudy (anonymous):

ok, thankyou

OpenStudy (anonymous):

No problem() good luck